a saturated solution of salt has been prepared at 298K . on completely evaporating 35 gram of solution 10g of salt was recovered . calulate the amount of salt that was dissolved in 100g of water while preparing the solution. what is the solubililty of salt at 298K

Grade:12th pass

1 Answers

Navin Rai
5 years ago
conc.  = (10g/35g)  x 100g = 28.57g salt in 100g soln.  
100 – 28.57= 71.43g water ;  ie. 28.57g salt in 71.43g water
In 100g water salt dissolved= (28.57/71.43) x 100= 39.997g salt
                                                                               ~40g salt